solve the system of equations using the addition method
7a+2b=13
8a-3b=-1
solution (?,?)

Answer by nerdybill(6959)

7a+2b=13
8a-3b=-1
.
Multiply the first equation by 3 and the second by 2 to get:
21a+6b=39
16a-6b=-2
.
Now, add the two equations together:
21a+6b=39
16a-6b=-2
---------
37a = 37
a = 1 (solution for a)
.
Substitute the above into the first equation and solve for b:
7a+2b=13
7(1)+2b=13
7+2b=13
2b = 4
b = 2 (solution for b)
.
So, (a,b) = (1,2)